OpenStudy (anonymous):

Brandon has a 55-gallon fish tank. He estimates he can drain the tank at an average rate of 11 gallons per minute. In a linear model of this situation, x represents the number of minutes the tank has been draining. Relative to this situation, what does the y-intercept represent? A. The y-intercept represents the rate at which the tank can be drained. B.The y-intercept represents the number of minutes it takes to drain the tank. C. The y-intercept represents the amount of water in the tank after it has been drained. D. The y-intercept represents the amount of water in the tank be

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

Okay, the equation I got was this: y = 55 -11x. That is y = 55gallons MINUS the amount of gallons drained at x amount of minutes. Make sense? the y-intercept is when x is equal to 0. So: y = 55 - (11 * 0) --> y = 55 What do you now know about the y-intercept?